Show your work.

1.

A. Using values for the class data that you computed in Part 2 of the project, construct a 99% confidence interval estimate for the true proportion of yellow candies using the class data as your sample. Remember that for this computation, n is the number of CANDIES for the entire class data. Include all your work, showing the formula used and appropriate values inserted (neatly written and scanned or typed) or including the appropriate calculator commands and inputs. (5 points)

B. Give an appropriate interpretation of your interval. (3 points)

C. Based on your interval for the true proportion of yellow candies, was the proportion of yellow candies in the single bag of candy you purchased a likely value for the true population proportion? Explain how you know using actual values from your data and computations. (5 points)

D. Using values you computed in Part 2 of the project, construct a 95% confidence interval estimate for the true mean number of candies per bag using the class data as your sample, but for this computation, n is the number of BAGS. Make sure you use the correct standard deviation from Part 2, which treats the class data set as a sample. Include all your work, showing the formula used and appropriate values inserted (neatly written and scanned or typed) or including the appropriate calculator commands and inputs. (5 points)

E. Give an appropriate interpretation of your interval. (3 points)

F. Based on your interval for the true mean number of candies per bag, was the total number of candies in the single bag you purchased a likely value for the population mean? Explain how you know using actual values from your data and computations. (5 points)

*part 2 project values:

green 1029

orange 1066

purple 969

red 1046

yellow 977

total: 5087

* Bag Purchased

green 14

orange 11

yellow 11

purple 7

red 13

total 56
Answer #1

a)

Formula: P hat + - z sqrt (p hat(1-p hat)/n

p^= X/n

X = 977 , n = 5087

z = 2.576

(   0.178 , 0.206 )

b)

With 99% confidence, the true proportion of yellow candies is between 0.178 and 0.206

c)

p^= 11/56 = 0.1964

The proportion of yellow candies is likely value 0.1964 is present in confidence interval

d)